OpenStudy (anonymous):

Question 1 Solve for x: |x| = 4 x = 0 and x = 4 x = 4 and x = −4 x = 4 x = −4 Question 2 Solve for x: −5|x + 1| = 1 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution Question 3 Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No Solution Question 4 Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution Question 5 Solve for x: |3x + 2| = 14 x = 4 and x = −4 x = 4 and x

OpenStudy (anonymous):

please explain it to me :)

OpenStudy (mathstudent55):

The form of absolute value equation you have is an expression with the variable inside the absolute value signs equals a number. When you have an absolute value equation of that form, the solution is obtained by setting up two equations separated by the word "or". The first equation is obtained by simply dropping the absolute value signs. The second equation is obtained by dropping the absolute value signs and taking the negative of what was inside the absolute signs. Then solve each equation and give both answers separated by the word "or". |x| = 4 becomes: x = 4 or -x = 4 x = 4 or x = -4

OpenStudy (mathstudent55):

Try 2.: -5|x + 1| = 1 First divide both sides by 5: |x + 1| = -1/5 Now it's in the form you need, but there's a problem. The absolute value of something cannot be a negative number, so here there is no solution.